# Benjamin Carleton

### v6

@@ 4 Results
Proof.    Assume $\mathcal{T}$ is one of our partition control types, that $\mathcal{E}$ is the election system of $\mathcal{T}$, that $W_{\mathcal{E}}\in\mathrm{P}$, and that (the decision problem) $\mathcal{T}$ is $\mathrm{NP}$-complete. $\mathrm{NP}$-hard: Let $f$ be any refinement of $\widehat{\Pi}_{\mathcal{T}}$. Since $\mathcal{T}$ is $\mathrm{NP}$-complete, it suffices to show $\mathcal{T}\in\mathrm{P}^{f}$ to show that $\mathrm{NP}\subseteq\mathrm{P}^{f}$. Our polynomial-time algorithm to decide $\mathcal{T}$ with function oracle $f$ proceeds as follows: If $x$ is not a triple of the form $(C,V,p)$ where $C$ is a set of candidates, $V$ is a vote collection over $C$ of the vote-type of $\mathcal{E}$, and $p\in C$, then reject. Otherwise, query the function oracle with $x$ and verify, in polynomial time (using the fact that $W_{\mathcal{E}}\in\mathrm{P}$), whether the oracle response is a solution to the control problem. If it is, accept, and otherwise reject.
$\mathrm{NP}$-easy: It is easy to see that there is a refinement of $\widehat{\Pi}_{\mathcal{T}}$, $h$, such that $h\in\mathrm{FP}^{\mathrm{NP}}$. Let us focus on $\mathcal{T}$ being a voter partition type. The candidate cases are exactly analogous except we build the candidate rather than the voter partition. On input $(C,V,p)$, $h$ makes sure that the input is of the form $(C,V,p)$, that $p\in C$, and that the votes in $V$ are of the type appropriate for $\mathcal{E}$; if not output $\perp$. Otherwise, use a single call to ${\rm SAT}$ to determine if $(C,V,p)\in\mathcal{T}$. If not, output $\perp$. Otherwise, by an easy binary search, with an $\mathrm{NP}$ oracle, we will construct a set $V_{1}$ such that $(V_{1},V-V_{1})$ is a solution to $\mathcal{T}$ for $(C,V,p)$. In fact, if we naturally encode partitions into binary strings, we can binary search, with any $\mathrm{NP}$-complete set such as ${\rm SAT}$ as our oracle, to find the lexicographically least encoding of a $V_{1}\subseteq V$ such that $(V_{1},V-V_{1})$ is a solution to $(C,V,p)$ with respect to $\mathcal{T}$, and then we will output $(V_{1},V-V_{1})$. (The “helper” $\mathrm{NP}$ set for the binary search is simply $\{(C,V,p,\mathcal{C})\mid(\exists\mathcal{C}^{\prime})[\mathcal{C}^{\prime}\geq_{\rm lex}\mathcal{C}$ and $\mathcal{C}^{\prime}$ encodes a $V_{1}$ such that $(V_{1},V-V_{1})$ is a solution to $(C,V,p)$ with respect to $\mathcal{T}]\}$. Since this is an $\mathrm{NP}$ set, questions to it can be polynomial-time transformed into questions to ${\rm SAT}$.)
Theorem 4.9 is a powerful tool that will help us determine the search complexity of many of our problems of interest as many of those decision problems have been shown to be $\mathrm{NP}$-complete in the literature. Unfortunately, much research on electoral control types has been in the unique-winner model, and so we need to establish the following new decision-complexity result before proceeding.
@@
Proof.    The analogous result in the unique-winner model was established by Hemaspaandra, Hemaspaandra, and Rothe [HHR07]. Our proof, which we include for completeness, closely follows their proof. We use the same construction (i.e., reduction), but our correctness argument involves small yet important modifications.
@@
We now state that construction that [HHR07] used for their NP-hardness reduction for Plurality-DC-PC-TP-UW, since we will use the same construction as our P-hardness reduction for Plurality-DC-PC-TP-NUW.
Construction 4.11 ([HHR07])
@@
      There are $42(m-k)+2n(k+1)+$ 4 votes of the form $c>w>\cdots$, where the $\cdots$” means that the remaining candidates are in some arbitrary order.
@@
      There are $52n(k+1)+$ 5 votes of the form $w>c>\cdots$.
@@
      For each $i\in\{1,\ldots,n\}$, there are $2(k+1)$ votes of the form $S_{i}>c>\cdots$, where “$S_{i}$” denotes the elements of $S_{i}$ in some arbitrary order.
@@
Proof of Claim 4.13.    Fix $D\subseteq B\cup\{w\}$ such that $c$ is not a winner of plurality election $(D\cup\{c\},V)$. We will show that the above three properties hold, by using a modified version of the argument used in [HHR07].
[…]
@@
Since $\ell$ is a nonnegative integer, the only value that it can have here is 0, and so it follows that $B^{\prime}$ is a hitting set of $S$ of size at most $k$.
@@
The “equivalently” follows from the fact that for every election system $\mathcal{E}$, -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ [HHM20].
@@
    $\mathrm{Plurality}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$
@@
Each of these four is NP-complete, the first by Proposition 4.10, the second is from [HHR07], and the remaining two are from Maushagen and Rothe [MR18]. The result follows from those four NP-completenesses, by Theorem 4.9.
@@
Table 1 provides, for the concrete election systems discussed in this paper, a summary of our results on the search complexity of each equivalence class of decision-collapsing control problems. This paper’s results establish that within each such equivalence class the search complexities must be polynomially related; the table is summarizing our results on whether those linked complexities are both “polynomial(-time computability)” or are both “SAT-equivalence.” (The leftmost column’s classification of which control types, for the three studied election systems, collapse as decision problems is due to Carleton et al. [${}^{+}$22], building on Hemaspaandra, Hemaspaandra, and Menton [HHM20].)
@@
Collapsing Decision Control Types Complexity of Their Search Versions
$\mathrm{Plurality}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Plurality}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ ${\rm SAT}$-equivalent
Corollaries 4.4 and 4.14
$\mathrm{Plurality}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Plurality}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Plurality}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Plurality}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ ${\rm SAT}$-equivalent
Corollaries 4.4 and 4.14
$\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$,$\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ polynomial
Corollary 4.5 and Theorem 4.8
$\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$,$\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ ${\rm SAT}$-equivalent
Corollaries 4.4 and 4.14
$\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ ${\rm SAT}$-equivalent
Corollaries 4.4 and 4.14
$\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ polynomial
Corollary 4.4 and Theorem 4.7
$\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ polynomial
Corollary 4.6 and Theorem 4.7
$\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$ polynomial
Corollary 4.5 and Theorem 4.8
$\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ polynomial
Corollary 4.6 and Theorem 4.7
$\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ polynomial
Corollary 4.5 and Theorem 4.7
$\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ polynomial
Corollary 4.5 and Theorem 4.8
@@
Table 1: For each collection of (decision-problem) collapsing partition control types for plurality, veto, and approval elections we have shown that their search problems are also of the same complexity. This table, for each, states whether the search problems are polynomial-time solvable (denoted “polynomial”) in the table or are ${\rm SAT}$-equivalent. Proposition 4.15 and Proposition 4.16 are not listed in the References column, since it is clear when they are being drawn on (see the discussion of this near the end of Section 4).
@@
$\mathrm{NP}$-easy: It suffices to show that $\widehat{\Pi}_{B}$ has a refinement in $\mathrm{FP}^{\rm SAT}$. Let $s$ denote a function witnessing $B\leq_{\textnormal{search}}^{p}A$. Since $A$ is ${\rm SAT}$-equivalent, there is a refinement $t$ of $\widehat{\Pi}_{A}$ such that $t\in\mathrm{FP}^{\rm SAT}$. But it also holds, for each $x$, that $t(x)$ is a solution for $x$ under $A$ if and only if ($t$ is not $\perp$ and) $s(x,t(x))$ is a solution for $x$ with respect to $B$. Thus the function, $h$, that on input $x$ is $\perp$ if $t(x)$ is $\perp$ and otherwise is $s(x,t(x))$ is a refinement of $\widehat{\Pi}_{B}$. But since $t\in\mathrm{FP}^{\rm SAT}$, clearly so also is $h$.
@@
Let $f$ be a polynomial-time computable refinement of $\widehat{\Pi}_{A}$ and let $g$ be a function witnessing $B\leq_{\textnormal{search}}^{p}A$. The function $h$ that on input $x$ is $\perp$ if $f(x)$ is $\perp$ and otherwise is $g(x,f(x))$ is clearly a polynomial-time computable refinement of $\widehat{\Pi}_{B}$.
@@
In this section, we have shown that for all known collapsing control-type pairs, even their search complexities are polynomially equivalent, given oracle access to the election system’s winner problem.

### v7

@@ 4.2 Concrete Search Complexities
Proof.  Assume $\mathcal{T}$ is one of our partition control types, that $\mathcal{E}$ is the election system of $\mathcal{T}$, that $W_{\mathcal{E}}\in\mathrm{P}$, and that (the decision problem) $\mathcal{T}$ is $\mathrm{NP}$-complete.
$\mathrm{NP}$-hard: Let $f$ be any refinement of $\widehat{\Pi}_{\mathcal{T}}$. Since $\mathcal{T}$ is $\mathrm{NP}$-complete, it suffices to show $\mathcal{T}\in\mathrm{P}^{f}$ to show that $\mathrm{NP}\subseteq\mathrm{P}^{f}$. Our polynomial-time algorithm to decide $\mathcal{T}$ with function oracle $f$ proceeds as follows: If $x$ is not a triple of the form $(C,V,p)$ where $C$ is a set of candidates, $V$ is a vote collection over $C$ of the vote-type of $\mathcal{E}$, and $p\in C$, then reject. Otherwise, query the function oracle with $x$ and verify, in polynomial time (using the fact that $W_{\mathcal{E}}\in\mathrm{P}$), whether the oracle response is a solution to the control problem. If it is, accept, and otherwise reject.
$\mathrm{NP}$-easy: It is easy to see that there is a refinement of $\widehat{\Pi}_{\mathcal{T}}$, $h$, such that $h\in\mathrm{FP}^{\mathrm{NP}}$. Let us focus on $\mathcal{T}$ being a voter partition type. The candidate cases are exactly analogous except we build the candidate rather than the voter partition. On input $(C,V,p)$, $h$ makes sure that the input is of the form $(C,V,p)$, that $p\in C$, and that the votes in $V$ are of the type appropriate for $\mathcal{E}$; if not output $\perp$. Otherwise, use a single call to $\rm SAT$} to determine if $(C,V,p)\in\mathcal{T}$. If not, output $\perp$. Otherwise, by an easy binary search, with an $\mathrm{NP}$ oracle, we will construct a set $V_{1}$ such that $(V_{1},V-V_{1})$ is a solution to $\mathcal{T}$ for $(C,V,p)$. In fact, if we naturally encode partitions into binary strings, we can binary search, with any $\mathrm{NP}$-complete set such as $\rm SAT$} as our oracle, to find the lexicographically least encoding of a $V_{1}\subseteq V$ such that $(V_{1},V-V_{1})$ is a solution to $(C,V,p)$ with respect to $\mathcal{T}$, and then we will output $(V_{1},V-V_{1})$. (The “helper” $\mathrm{NP}$ set for the binary search is simply $\{(C,V,p,\mathcal{C})\mid(\exists\mathcal{C}^{\prime})[\mathcal{C}^{\prime}\geq_{\rm lex}\mathcal{C}$ and $\mathcal{C}^{\prime}$ encodes a $V_{1}$ such that $(V_{1},V-V_{1})$ is a solution to $(C,V,p)$ with respect to $\mathcal{T}]\}$. Since this is an $\mathrm{NP}$ set, questions to it can be polynomial-time transformed into questions to $\rm SAT$}.) ❑
Theorem 4.9 is a tool that will help us determine the search complexity of many of our problems of interest, since many of those decision problems have been shown to be $\mathrm{NP}$-complete in the literature. Unfortunately, much research on electoral control types has been in the unique-winner model, and so we need to establish the following new decision-complexity result before proceeding.
@@
Proof.  The analogous result in the unique-winner model was established by Hemaspaandra, Hemaspaandra, and Rothe [HHR07a]. Our proof, which we include for completeness, closely follows their proof. We use the same construction (i.e., reduction), but our correctness argument involves small yet important modifications.
@@
We now state the construction that [HHR07a] used for their NP-hardness reduction for Plurality-DC-PC-TP-UW, since we will use the same construction as our NP-hardness reduction for Plurality-DC-PC-TP-NUW.
Construction 4.11 ([HHR07a])
@@
      There are 2(mk)+2n(k+1)+42(m-k)+2n(k+1)+4 votes of the form $c>w>\cdots$, where the $\cdots$” means that the remaining candidates are in some arbitrary order.
@@
      There are $2n(k+1)+5$ votes of the form $w>c>\cdots$.
@@
      For each $i\in\{1,\ldots,n\}$, there are $2(k+1)$ votes of the form $S_{i}>c>\cdots$, where $S_{i}$” denotes the elements of $S_{i}$ in some arbitrary order.
@@
Proof of Claim 4.13.  Fix $D\subseteq B\cup\{w\}$ such that $c$ is not a winner of plurality election $(D\cup\{c\},V)$. We will show that the above three properties hold, by using a modified version of the argument used in [HHR07a].
For a given election, we will let $\mathit{score}(d)$ denote the number of votes that rank candidate $d$ first in that election.
First, notice that for each $b\in D\cap B$, $\mathit{score}(b)<score(c)$ in $(D\cup\{c\},V)$. Since $c$ is not a winner of that election, it must hold that $w$ is the unique winner of that election, and thus $\mathit{score}(w)>\mathit{score}(c)$.
@@
Since $\ell$ is a nonnegative integer, the only value that it can have here is 0, and so it follows that $B^{\prime}$ is a hitting set of $S$ of size at most $k$. ❑
@@
The “equivalently” follows from the fact that for every election system $\mathcal{E}$, -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}=\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ [HHM20]. ❑
@@
    $\mathrm{Plurality}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$
@@
Each of these four is NP-complete, the first by Proposition 4.10, the second is from [HHR07a], and the remaining two are from Maushagen and Rothe [MR18]. The result follows from those four NP-completenesses, by Theorem 4.9. ❑
@@
Table 1 provides, for the concrete election systems discussed in this paper, a summary of our results on the search complexity of each equivalence class of decision-collapsing control problems. This paper’s results establish that within each such equivalence class the search complexities must be polynomially related; the table is summarizing our results on whether those linked complexities are both “polynomial(-time computability)” or are both “SAT-equivalence.” (The leftmost column’s classification of which control types, for the three studied election systems, collapse as decision problems is due to Carleton et al. [CCH+24], building on Hemaspaandra, Hemaspaandra, and Menton [HHM20].)
@@
Collapsing Decision Control Types Complexity of Their Search Versions References
$\mathrm{Plurality}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Plurality}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ $\rm SAT$}-equivalent Corollaries 4.4 and 4.14
$\mathrm{Plurality}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Plurality}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Plurality}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Plurality}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ $\rm SAT$}-equivalent Corollaries 4.4 and 4.14
$\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$$\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ polynomial Corollary 4.5 and Theorem 4.8
$\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$$\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ $\rm SAT$}-equivalent Corollaries 4.4 and 4.14
$\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ $\rm SAT$}-equivalent Corollaries 4.4 and 4.14
$\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ polynomial Corollary 4.4 and Theorem 4.7
$\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ polynomial Corollary 4.6 and Theorem 4.7
$\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$ polynomial Corollary 4.5 and Theorem 4.8
$\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ polynomial Corollary 4.6 and Theorem 4.7
$\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{NUW}$ polynomial Corollary 4.5 and Theorem 4.7
$\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, $\mathrm{Approval}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\allowbreak\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ polynomial Corollary 4.5 and Theorem 4.8
@@
Table 1: For each collection of (decision-problem) collapsing partition control types for plurality, veto, and approval elections, we have shown that their search problems are also of the same complexity. This table, for each, states whether the search problems are polynomial-time solvable (denoted “polynomial” in the table) or are $\rm SAT$}-equivalent. Proposition 4.15 and Proposition 4.16 are not listed in the References column, since it is clear when they are being drawn on (see the discussion of this near the end of Section 4).
@@
$\mathrm{NP}$-easy: It suffices to show that $\widehat{\Pi}_{B}$ has a refinement in $\mathrm{FP}^{\rm SAT}$. Let $s$ denote a function witnessing $B\leq_{\textnormal{search}}^{p}A$. Since $A$ is $\rm SAT$}-equivalent, there is a refinement $t$ of $\widehat{\Pi}_{A}$ such that $t\in\mathrm{FP}^{\rm SAT}$. But it also holds, for each $x$, that $t(x)$ is a solution for $x$ under $A$ if and only if ($t$ is not $\perp$ and) $s(x,t(x))$ is a solution for $x$ with respect to $B$. Thus the function, $h$, that on input $x$ is $\perp$ if $t(x)$ is $\perp$ and otherwise is $s(x,t(x))$ is a refinement of $\widehat{\Pi}_{B}$. But since $t\in\mathrm{FP}^{\rm SAT}$, clearly so also is $h$. ❑
@@
Let $f$ be a polynomial-time computable refinement of $\widehat{\Pi}_{A}$ and let $g$ be a function witnessing $B\leq_{\textnormal{search}}^{p}A$. The function $h$ that on input $x$ is $\perp$ if $f(x)$ is $\perp$ and otherwise is $g(x,f(x))$ is clearly a polynomial-time computable refinement of $\widehat{\Pi}_{B}$. ❑
@@
Section 4 has now achieved the broad set of results that its first paragraph promised, and as will also be summarized in the Conclusion section.